# Mass and energy cascade in collisionless dark matter flow and relevant constraints on the dark matter particle massA.6

### v3

@@ 4.2 Formulating the mass cascade in mass space
Figure 4 presents the evolution of the mass flux function $\Pi_{m}\left(m,z\right)$ from the Virgo simulation. Similarly, a scale-independent mass flux $\varepsilon_{m}(z)$ can be clearly identified for the halo mass $m_{h}<m_{h}^{*}$. The propagation range with scale-independent mass flux $\varepsilon_{m}$ is formed around $z=0.3$ and gradually expands in mass space. The time to reach the statistical state with a scale-independent rate of cascade is dependent on the simulation resolution. The higher mass resolution in the Illustris simulation leads to an earlier formation of the propagation range. The simulation with a higher resolution establishes the statistically steady state at a higher redshift.
Figure 4: The variation of mass flux $-\Pi_{m}\left(m_{h},z\right)$ from the Virgo simulation using halo group mass $m_{g}\left(a\right)$ at two different redshifts z (Eq. (16)). A scale-independent mass flux $\varepsilon_{m}\left(a\right)$ can be found for haloes smaller than the characteristic mass $m_{h}<m_{h}^{*}$. The negative mass flux $\varepsilon_{m}(z)<0$ indicates an inverse mass cascade from small to large scales. A mass propagation range with a scale-independent $\varepsilon_{m})$ is formed at around $z=0.3$, much later than the Illustris simulation in Fig 3 due to a lower mass resolution.
4.3 Waiting time $\tau_{g}$ and universal scaling laws
The mass cascade involves a waiting time $\tau_{g}$ for haloes to merge and migrate from one mass scale to the neighboring mass scale (Eq. (10)). Combining $\tau_{g}$ with the scale-independent rate of the mass cascade $\varepsilon_{m}$, some important universal scaling laws can be easily identified.
The waiting time $\tau_{g}$ is inversely proportional to the halo potential $\tau_{g}\propto\Phi_{h}^{-1}=(Gm_{h}/r_{h})^{-1}$ (Xu, 2023b), where $r_{h}$ is the size of that halo. This is also explained in Section 5.2 (Eq. (55)), where the halo merging and accretion can be related to the particle migration in haloes. Larger haloes with a higher gravitational potential tend to accumulate mass more quickly with a shorter waiting time $\tau_{g}$. Putting all of these together, we can write
$\begin{split}&\frac{dm_{h}}{dt}=\frac{m_{p}}{\tau_{g}},\\ &\tau_{g}\propto\Phi_{h}^{-1}\propto(Gm_{h}/r_{h})^{-1},\\ &\rho_{h}=\frac{m_{h}}{\frac{4}{3}\pi{r_{h}}^{3}}=\Delta_{c}\rho_{DM}\propto a^{-3},\end{split}$ (18)
where $m_{p}$ is the mass of a single merger (or dark matter particle). The first equation describes that, on average, the halo accretes a mass of $m_{p}$ during every waiting time $\tau_{g}$. This equation describes the average mass accretion. The second equation quantifies the waiting time. The third equation relates the average halo density to the mean density $\rho_{DM}$, where $\Delta_{c}=18\pi^{2}$ is the critical density ratio. Solving these equations leads to:
Therefore, for a halo with a given mass $m_{h}$, the waiting time can be written as $\tau_{g}\propto a{m_{h}^{-\lambda}}$, where $\lambda$ is a key halo parameter and $\lambda=2/3$ is expected for large haloes.
From the definitions of the rate of mass cascade $\varepsilon_{m}$ and the halo group mass $m_{g}$ (Eq. (10)), we have
$\begin{split}&-\varepsilon_{m}(z)=\frac{m_{g}(m_{h})}{\tau_{g}(a,m_{h})}=\frac{d}{dt}M_{h}(t)\\ &m_{g}(m_{h})=M_{h}(t)f_{M}(m_{h},z)m_{p}=n_{h}m_{h}.\end{split}$ (20)
Here, $n_{h}$ is the number of haloes with the same mass $m_{h}$. The mass of the halo group $m_{g}$ is independent of the redshift $z$ due to the scale-independent mass cascade. That is, the mass flux into a halo group is equal to the mass flux out of the same group, so the mass of the group $m_{g}$ is a function of $m_{h}$ only and is constant over time. In addition, with mass injected from the smallest scale $m_{p}$ at a scale-independent rate of $\varepsilon_{m}$, the rate of the mass cascade also equals the rate of change in the total mass $M_{h}(t)$ in all haloes.
Figure 5: The variation of halo group mass $m_{g}(m_{h})$ for all haloes of the same mass $m_{h}$ at different redshifts $z$ from Illustris-1-Dark simulation. The figure demonstrates the small-scale permanence of the group mass $m_{g}$. Once the statistically steady state is established ($z\leq 4$), the rate of the inverse mass cascade $\varepsilon_{m}$ becomes scale independent. The mass flux into a halo group balances the mass flux out of the same group such that the group mass $m_{g}$ at different redshifts $z$ collapses into a time-independent power law $m_{g}\propto m_{h}^{-\lambda}$ on small mass scales. This is the so-called "propagation range," where the mass cascaded from the smallest (DM particle) scale is propagated to larger scales, that is, a constant group mass $m_{g}$. The cascaded mass is eventually consumed to grow haloes greater than $m_{h}^{*}$ (the "deposition range"). The propagation range gradually extends to large scales ($m_{h}^{*}(a)$ increases over time) due to the continuous inverse mass cascade.
Combining Eq. (20) with Eq. (19) for $\tau_{g}$, universal scaling laws can be identified when the system reaches the statistically steady state with a scale-independent rate $\varepsilon_{m}$ for mass cascade. First,
The halo group mass follows a simple power law $m_{g}\propto m_{h}^{-\lambda}$ that is redshift independent. This is the so-called small-scale permanence demonstrated in Fig. 5, i.e., the group mass $m_{g}(m_{h})$ at different redshift $z$ collapses into the same scaling law $m_{g}\propto m_{h}^{-\lambda}$.
Using Eq. (20), the halo mass function $f_{M}$ satisfies
$\begin{split}&f_{M}(m_{h},z)=\frac{m_{g}(m_{h})}{M_{h}(t)m_{p}}\propto{m_{h}}^{-2/3}a^{-1/2},\end{split}$ (22)
where the halo mass function follows a simple power law in propagation range $m_{h}<m_{h}^{*}$. By absorbing the redshift dependence into the characteristic mass $m_{h}^{*}$, we may express the halo mass function in a very simple power-law form (also see Eq. (30)). Combined with Eq. (22), the scaling for the evolution of $m_{h}^{*}$ can be obtained,
$\begin{split}&f_{M}(m_{h}^{*},m_{h})\propto\frac{1}{m_{h}^{*}}\left(\frac{m_{h}}{m_{h}^{*}}\right)^{-\lambda}\quad\textrm{leads to}\quad m_{h}^{*}\propto a^{\frac{1}{2(1-\lambda)}}.\end{split}$ (23)
Finally, using Eqs. (19) ($\lambda=2/3$) and Eq. (23), for haloes with the characteristic mass $m_{h}^{*}$ and size $r_{h}^{*}$, we found the scaling
$r_{h}^{*}\propto t,\quad m_{h}^{*}\propto t\quad\textrm{and}\quad\tau_{g}^{*}\propto t^{0}$ (24)
in the matter-dominant era. The waiting time $\tau_{g}$ for haloes with characteristic mass $m_{h}^{*}$ is constant over time in the matter era. The total mass in all haloes $M_{h}\propto a^{1/2}\propto t^{1/3}$ (Eq. (21)).
Figure 6: The evolution of the characteristic halo mass $m_{h}^{*}(t)$ (red dashed) and total halo mass $M_{h}$ (blue dashed) in the matter era, where $m_{h}^{*}\propto a^{3/2}\propto t$ and $M_{h}\propto a^{1/2}\propto t^{1/3}$. Results from Illustris simulations (symbols) are also presented that are in agreement with the prediction of Eqs. (21) and (24). The deviation at high redshift can be due to the limited mass resolution.
Figure 6 presents the time variation of $m_{h}^{*}$ and $M_{h}$ from the Illustris simulations. After the system reaches a statistically steady state with a scale-independent rate of cascade, the characteristic mass $m_{h}^{*}\propto a^{3/2}$ and the total halo mass $M_{h}\propto a^{1/2}$ confirm the scaling laws derived from the mass cascade (Eqs. (21) and (23)).
4.4 Halo random walk and halo mass function
The power law mass function in the propagation range can be derived from universal scaling laws (Eq. (22)), while the complete halo mass function in the entire range of mass scales can be derived from the random walk of haloes in mass space.
First, the inverse mass cascade leads to a random walk of haloes in the mass space that mimics the random work of particles for diffusion. As shown in Fig. 2, haloes continuously migrate in mass space from one scale ($m_{h}$) to the neighboring scale ($m_{h}+m$) by merging with single mergers of mass $m$. There is a waiting time $\tau_{g}$ associated with the halo random walk. The waiting is position-dependent, i.e., dependent on the position $m_{h}$ in mass space (Eq. (19)).
Second, the particle random walk for diffusion leads to an evolving particle distribution according to the Fokker-Planck equation. Similarly, the random walk of haloes in mass space leads to a probability distribution to find a halo at a given mass. The distribution of haloes with respect to the halo mass, i.e., the halo mass function, is naturally obtained from the corresponding Fokker-Planck equation.
@@ 4.4 Halo random walk and halo mass function
This approach is mathematically convenient and reveals some fundamental aspects of the halo mass function. The evolving halo mass function continuously maximizes the system entropy contained in the distribution of the halo mass, just as the evolving particle distribution maximizes the entropy contained in the particle distribution. Mathematically, the random walk of haloes in mass space can be described by the stochastic Langevin equation (Xu, 2023b)
@@
Unlike the standard particle random walk for diffusion, the waiting time $\tau_{g}$ for the halo random walk depends on the position in the mass space (or the halo mass $m_{h}$). For a power-law waiting time $\tau_{g}\propto{m_{h}^{-\lambda}}$ ($\lambda=2/3$ from Eq. (19)), the position-dependent diffusivity should take the form of (comparing Eq. (25) to Eq. (18))
@@
Here, $D_{p0}(t)$ is a proportional constant for diffusivity $D_{p}$. The white Gaussian noise $\varsigma(t)$ satisfies the covariance $\langle\varsigma(t)\varsigma(t^{{}^{\prime}})\rangle=\delta(t-t^{{}^{\prime}})$ with zero mean $\langle\varsigma(t)\rangle=0$.
Next, in Stratonovich’s interpretation (Stratonovich, 1966; Metzler et al., 2014), the Langevin equation (Eq. (25)) yields the Fokker-Planck equation for the halo mass function $f_{M}(m_{h},t)$ (in analogue to the particle diffusion):
@@
The characteristic mass $m_{h}^{*}(t)$ can be related to the mean square displacement in mass space as
[…]
where $\gamma_{0}$ is a proportional constant. With $\lambda<1$ and the exponent ${1/\left(1-\lambda\right)}\geq 1$ in Eq. (29), the random walk of haloes in mass space is essentially a super-diffusion.
Now, the halo mass function can be rewritten as
$f_{M}\left(m_{h},t\right)=\frac{\left(1-\lambda\right)}{m_{h}^{*}\sqrt{\pi\eta_{0}}}\left(\frac{m_{h}^{*}}{m_{h}}\right)^{\lambda}\exp\left[-\frac{1}{4\eta_{0}}\left(\frac{m_{h}}{m_{h}^{*}}\right)^{2-2\lambda}\right],$ (30)
@@
The time dependence of $f_{M}$ is absorbed into $m_{h}^{*}$. Clearly, Eq. (30) reduces to the Press-Schechter (PS) mass function if $\lambda=2/3$ and $\eta_{0}={1/2}$ or $\gamma_{0}=15$. Note that this approach for deriving the halo mass function does not rely on any collapse model (either spherical or ellipsoidal). By introducing the dimensionless variable $\nu=(m_{h}/m_{h}^{*})^{2/3}$, mass function can be expressed in terms of $\nu$
@@
where parameter $p=3\left(1-\lambda\right)$ is on the order of unity.
Figure 7: Comparison between different mass functions (log(f($\nu$))) in Appendix A.6 and cosmological simulations. The black square represents the mass function from Virgo simulation at z=0. The color symbols represent the mass function from the Illustris simulation at different redshifts $z$. The PS mass function underestimates the mass in large haloes. The fitted JK mass function matches the simulation for a given range of halo size but not the entire range. The double-$\lambda$ mass function (Eq. (35)) matches both the simulation and ST mass function for the entire range.
In principle, two different $\lambda$ (that is, double-$\lambda$) are required for two different ranges (propagation and deposition range). The halo mass function in Eq. (30) can be naturally generalized to a double-$\lambda$ halo mass function with $\lambda_{1}$ and $\lambda_{2}$ for propagation and deposition ranges, respectively (Xu, 2023b). Based on this idea, a double-$\lambda$ mass function can be proposed based on the single $\lambda$ mass function in Eq. (30),
@@
where $q={\left(1-\lambda_{1}\right)/\left(1-\lambda_{2}\right)}$ is the ratio between two $\lambda$ values to satisfy the normalization constraint. The kth order moment of the double-$\lambda$ mass function is
@@
With the dimensionless variable defined as $v=\left({m_{h}/m_{h}^{*}}\right)^{2-2\lambda_{2}}$, the dimensionless double-$\lambda$ mass function reads,

### v4

@@ 2 Universal transition spectrum with n=-1
The solid lines in Fig. 3 plot three different ratio parameters $\alpha_{x}$=1/2, 2/3, and 3/4. Peebles’ prediction corresponds to the special case with $\alpha_{x}=0$ or $k_{NL1}=k_{NL2}$ (dashed line). Compared with simulation data, an expanding universal transition range with a ratio parameter $\alpha_{x}=1/2$ is preferred. This ratio can be related to the characteristic halo mass $m_{h}^{*}\propto a^{3/2}$ (or a moving size $r_{hc}^{*}\propto a^{1/2}$) from the theory of the mass cascade (Eq. (32)). The increase in characteristic halo size $r_{hc}$ reflects the expanding universal range and ratio $\alpha_{x}$.
This section discusses the existence of a universal transition range and its effects on the asymptotic spectrum index $n_{a}$. That transition range admits a unique and universal power-law spectrum $P(k)\propto k^{-1}$ that expands with time as $\propto a^{1/2}$. In the next, a cascade theory is developed to interpret the formation and evolution of such a universal spectrum, which also confirms the same spectrum index of 1-1 (Eq. (31) in Section 3.2). Similarly, associated with the mass and energy cascade, universal scaling laws for halo mass functions and density profiles also exist near the scale $m_{h}^{*}$ (Sections 3.2 and 4.2).
3 Mass cascade and halo mass function
The universal spectrum suggests a statistically steady state established in a nonequilibrium system that emerges from some "cascade" process, largely losing the memory of initial conditions. In this section, we introduce the notion of a mass and energy "cascade" across dark matter haloes spanning a range of masses. Because the dark component is effectively collisionless, it cannot sustain a classical turbulent cascade mediated by viscosity and vorticity. We therefore employ “cascade” in a generalized sense to denote the redistribution of mass and energy across halo-mass scales that drives the system toward its asymptotic equilibrium, rather than a Kolmogorov-type viscous turbulent process. A salient aspect of this evolution is the emergence of a statistically stationary state characterized by a scale-invariant flux of mass and energy through the halo mass space.
3.1 Halo random walk and mass function
The distribution of haloes depends both on the initial conditions and on gravitational dynamics. The Press-Schechter formalism suggests a mass function $\propto{m_{h}}^{\frac{n-3}{6}}$ dependent only on the spectrum index $n$ (Eq. (24)). Using Eq. (14) for the asymptotic spectrum index $n_{a}$, the asymptotic slope $\lambda_{a}$ of the mass function $m_{h}^{-\lambda_{a}}$ can be related to the initial power-law spectrum index $n_{i}$ as
[…]
In hierarchical (bottom-up) structure formation, low-mass haloes collapse at high redshift from peaks of the primordial density field. Their distributions are therefore largely determined by the properties of these peaks and retain a strong imprint of the primordial power spectrum. In contrast, high-mass haloes assemble later through repeated mergers and smooth accretion. This prolonged and violent assembly history progressively erases information about the initial conditions, driving the system toward a quasi-universal state. Consequently, the distribution for haloes around characteristic mass $m_{h}^{*}$ is expected to be universal, governed primarily by gravitational dynamics, and less sensitive to the primordial spectrum.
To clarify this behavior, we consider an idealized initial condition in which the halo population is sharply concentrated at $m_{h}$=0 in mass space, representing the initial simultaneous collapse of an extremely large number of small and high-density proto-haloes at a high redshift and at various locations throughout the universe. In this limit, the influence of the primordial spectrum is confined to a small region near $m_{h}=0$, whereas subsequent evolution proceeds subject to self-gravity in an expanding dark matter background. We adopted an effective mean-field description to describe structure formation via smooth mass accretion from the background (Fig. 4). The evolution of the halo distribution in mass space is modeled as a continuous-time random walk with a mass-dependent waiting time. In the gravity-dominated regime, this waiting time is controlled by the halo potential $\tau_{g}(m_{h},z)\propto\Phi(m_{h})^{-1}\propto m_{h}^{-\lambda}$ (Eq. (75)). The larger the haloes, the deeper the halo potential, the hotter the local temperature, the shorter the residence times, and the faster the mass accretion. When gravity dominates the merging rate, a universal $\lambda=2/3$ can be obtained, as shown in Eq. (31). If the primordial spectrum dominates, the exponent $\lambda$ is expected to be related to the index of the initial spectrum, or $\lambda=(3-n)/6$, as shown in Eq. (25).
Figure 4: Schematic plot for the halo random walk and mass and energy cascade in halo mass space. Haloes of mass $M$ merging with a single merger (free particles of mass $m$) leads to a mass flux to a larger scale $M+m$ (smooth accretion), that is, the halo of mass $M$ walking into the next mass scale $M+m$ after merging. For a given halo of mass $m_{h}$, a merge occurs with an average waiting time $\tau_{g}(m_{h},z)$. Because haloes have finite mass and kinetic energy, continuous merging facilitates an inverse mass and kinetic energy cascade from small to large scales and a direct cascade of potential energy from large to small scales. The scale-independent mass and energy flux ($\varepsilon_{m}$ and $\varepsilon_{u}$ are independent of mass scale $m_{h}$) are expected in the propagation range ($m=m_{p}<m_{h}<m_{h}^{*}$), a typical feature of a statistical steady state in a nonequilibrium system. The rate of cascade becomes scale-dependent in the deposition range ($m_{h}>m_{h}^{*}$) (see Figs. 7, 8 for $\varepsilon_{m}$ and Fig. 10 for $\varepsilon_{u}$).
Figure 4 describes the random walk of haloes in mass space. The merging of a halo (mass $M$) with a single mass merger $m$ (smooth accretion, depending on the mass resolution in the N-body simulations) leads to a continuous mass and energy flux to a larger mass scale $M+m$, that is, the halo of mass $M$ walking into the next mass scale $M+m$ after merging. For a given halo mass $m_{h}$, the merge event occurs with an average waiting time $\tau_{g}(m_{h},z)$. Therefore, haloes continuously migrate in mass space from one scale ($m_{h}$) to the neighboring scale ($m_{h}+m$), enabled by the merging with single mergers (smooth accretion). A diffusion-like particle random walk results in an evolving particle distribution, as described by the Fokker-Planck equation. Similarly, the random walk of haloes in mass space leads to a probability distribution to find a halo at a given mass. The distribution of haloes with respect to the halo mass, i.e., the halo mass function, can be naturally obtained from the corresponding Fokker-Planck equation. This approach for the mass function does not rely on any collapse model (either spherical or ellipsoidal).
@@ 3.1 Halo random walk and mass function
This approach is mathematically convenient, fundamentally different from the Press-Schechter formalism, and reveals important aspects of the halo mass function. The evolving halo mass function continuously maximizes the system entropy contained in the distribution of the halo mass, just as the evolving particle distribution maximizes the entropy contained in the particle distribution. Mathematically, the random walk of haloes in mass space can be described by a stochastic Langevin equation (Xu:2023-Dark-matter-halo-mass-functions-and)
@@
Unlike the standard particle random walk for diffusion, the waiting time $\tau_{g}$ for the halo random walk is scale-dependent. For a power-law waiting time $\tau_{g}\propto{m_{h}^{-\lambda}}$, the scale-dependent diffusivity reads
@@
Here, $D_{p0}(t)$ is a proportional constant for the diffusivity $D_{p}$. The white Gaussian noise $\varsigma(t)$ satisfies the covariance $\langle\varsigma(t)\varsigma(t^{{}^{\prime}})\rangle=\delta(t-t^{{}^{\prime}})$ with zero mean $\langle\varsigma(t)\rangle=0$.
Next, in Stratonovich’s interpretation (Stratonovich:1966-A-new-representation-for-stoch; Metzler:2014-Anomalous-diffusion-models-and), the Langevin equation (Eq. (16)) yields the Fokker-Planck equation for the halo mass function $f_{M}(m_{h},t)$:
@@
The characteristic mass $m_{h}^{*}(t)$ can be determined numerically from N-body simulations (Xu:2023-Dark-matter-halo-mass-functions-and) (Fig. 6). In principle, it is related to the mean square displacement in the mass space as
$\begin{split}&\left\langle m_{h}^{2}\right\rangle=\int_{0}^{\infty}f_{M}\left(m_{h},t\right)m_{h}^{2}dm_{h}\\ &=\frac{1}{\sqrt{\pi}}\Gamma\left(\frac{3-\lambda}{2-2\lambda}\right)\left[4\left(1-\lambda\right)^{2}D_{p0}t\right]^{\frac{1}{1-\lambda}}\equiv\gamma_{0}m_{h}^{*2}.\end{split}$ (20)
where $\gamma_{0}$ is a proportional constant. With $\lambda<1$ and the exponent ${1/\left(1-\lambda\right)}\geq 1$ in Eq. (20), the random walk of haloes in mass space is essentially a super-diffusion. With $m_{h}^{*}$ as the characteristic mass scale, the halo mass function can be rewritten as
$f_{M}\left(m_{h},t\right)=\frac{\left(1-\lambda\right)}{\sqrt{\pi\eta_{0}}m_{h}^{*}}\left(\frac{m_{h}}{m_{h}^{*}}\right)^{-\lambda}\exp\left[-\frac{1}{4\eta_{0}}\left(\frac{m_{h}}{m_{h}^{*}}\right)^{2-2\lambda}\right],$ (21)
@@
From this derivation, the parameter $\lambda$ for waiting time $\tau_{g}\propto m_{h}^{-\lambda}$ is essentially the slope of the halo mass function. By introducing a dimensionless variable $\nu=(m_{h}/m_{h}^{*})^{2/3}$, the mass function can be expressed in terms of $\nu$
@@
where the parameter $p=3\left(1-\lambda\right)$ is on the order of unity.
The Press-Schechter (PS) mass function shares a similar form, but involves a spectrum index $n$
$\begin{split}f_{PS}\left(m_{h}\right)=\frac{3+n}{3\sqrt{2\pi}}\frac{1}{M^{*}}\left(\frac{m_{h}}{M^{*}}\right)^{-\frac{3-n}{6}}\exp\left[-\frac{1}{2}\left(\frac{m_{h}}{M^{*}}\right)^{\frac{3+n}{3}}\right].\end{split}$ (24)
When two mass functions are compared (Eqs. (21) and (24)), the parameter $\lambda$ is related to the spectrum index $n$
$\lambda=\frac{3-n}{6}.$ (25)
Clearly, our mass function reduces to the PS mass function if $\lambda=(3-n)/6$ and $\eta_{0}={1/2}$ or $\gamma_{0}=15$. However, without considering the universal transition range, the mass scale $M^{*}$ in the Press-Schechter mass function is simply linked to the critical scale $k_{NL1}$ with $M^{*}\propto a^{\frac{6}{n+3}}$ in linear theory. While our characteristic mass $m_{h}^{*}$ initially tracks $M^{*}$, it becomes smaller than $M^{*}$, and scales as $m_{h}^{*}\propto a^{3/2}$ as the universal transition range develops and broadens (Fig. 1).
The parameter $\lambda$ is the slope of the mass function, which also quantifies the waiting time or the merging rate. A more negative $n$ leads to a larger $\lambda$ and less frequent merging. For the CDM spectrum, $\lambda$ could increase and approach 1 with decreasing mass scales (or increasing wavenumber k). Due to the dominance of initial power spectrum and gravitational dynamics on different scales, the simplest general scenario is to have two different "effective" $\lambda$ (that is, a double-$\lambda$ model) for two different ranges. The fully nonlinear region on small scales ($m_{h}\ll m_{h}^{*}$) retains a strong memory of initial conditions, whereas the universal transition range ($m_{h}\sim m_{h}^{*}$) is governed by gravitational dynamics. The halo mass function in Eq. (21) can be naturally generalized to include two $\lambda$ parameters, with $\lambda_{1}$ for the effects of initial spectrum and $\lambda_{2}$ for gravity-dominated merging (Xu:2023-Dark-matter-halo-mass-functions-and). Based on this idea, a double-$\lambda$ mass function can be naturally constructed from Eq. (21),
@@
where $q={\left(1-\lambda_{1}\right)/\left(1-\lambda_{2}\right)}$ is the ratio between two $\lambda$ values to satisfy the normalization. The kth order moment of the double-$\lambda$ mass function is
@@
With a dimensionless variable defined as $v=\left({m_{h}/m_{h}^{*}}\right)^{2-2\lambda_{2}}$, the dimensionless double-$\lambda$ mass function reads
@@
Figure 5: Comparison between different mass functions (log(f($\nu$))) in Appendix C.6 and cosmological simulations. The black square represents the mass function from the Virgo SCDM simulation at z=0. The color symbols represent the mass function from the Illustris simulation at different redshifts $z$. The PS mass function underestimates the mass in large haloes. The fitted JK mass function matches the simulation for a given range of halo size, but not the entire range. The double-$\lambda$ mass function (Eq. (28)) matches both the simulation and ST mass function for the entire range.